OpenStudy (anonymous):

What is the value of X in 3/5 * X = 12? *

OpenStudy (alexandervonhumboldt2):

so easy" \[\frac{ 3 }{ 5 }x=12\]

OpenStudy (anonymous):

Thankyou :) can you explain a little more

OpenStudy (alexandervonhumboldt2):

cross multply: \[x=12 * \frac{ 5 }{ 3 }=\frac{ 12*5 }{ 3 }=?\]

OpenStudy (michele_laino):

if I multiply both sides by 5, I get: \[\frac{3}{5}x \cdot 5 = 12 \cdot 5\]

OpenStudy (anonymous):

:D i got 60 michele

OpenStudy (michele_laino):

please simpify

OpenStudy (alexandervonhumboldt2):

yeah now divide by 3

OpenStudy (anonymous):

i got 60 from 12 X 5

OpenStudy (michele_laino):

more precisely, we get: \[3x = 60\] since \[\frac{3}{5} \cdot 5 = 3\]

OpenStudy (anonymous):

:) thankyou this is awsome

OpenStudy (michele_laino):

then if I divide both sides by \(3\), I get: \[\frac{{3x}}{3} = \frac{{60}}{3}\] again, please simplify

OpenStudy (anonymous):

okay 1 moment michele

OpenStudy (anonymous):

20 is 60/3

OpenStudy (michele_laino):

that's right! since \[\frac{3}{3} = 1\] then we can write: \(x=20\)

OpenStudy (anonymous):

and 3/3 is 1

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

so the anser is ...

OpenStudy (anonymous):

x = 20

OpenStudy (michele_laino):

that's right!!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

YAY! u are awsome

OpenStudy (alexandervonhumboldt2):

yeh it is 20

OpenStudy (michele_laino):

thanks!! :D

OpenStudy (anonymous):

thx now im going to my next questin

OpenStudy (alexandervonhumboldt2):

oke

OpenStudy (michele_laino):

ok!

imqwerty (imqwerty):

don't forget to rate the QHs :)